OpenStudy (anonymous):

the square root of 3x-14 and -8=0

OpenStudy (anonymous):

foind the zeros

OpenStudy (anonymous):

\[\sqrt{3x-14}=8\] \[3x-14=64\] \[3x=78\] \[x=78\div3=26\]

OpenStudy (nilankshi):

u ar telling u ar not genius sww u ar

OpenStudy (nilankshi):

see

OpenStudy (anonymous):

what?